In a complex assembly operation, it is found that the learning curve rate is 70%. The standard time of 3 minutes per assembly is reached after the 110th unit. Find

i. The time required for the very first unit.
ii. The time required for the 200th unit.
Answer #1

The standard formula,

Time for the nth unit produced = Time for 1st unit produced*no of units produced^(ln of learning percentage as a decimal/ln2)

i.

3 = Time for 1st unit produced*110^(LN(0.7)/LN(2))

Time for 1st unit produced = 3/(110^(LN(0.7)/LN(2))) = 33.69

ii.

Time for the 200th unit produced = 33.69*(200^(LN(0.7)/LN(2))) = 2.21 minutes